OpenStudy (megannicole51):

A pile of earth standing on flat ground by an abandoned mine has height 15 meters. The ground is the xy-plane; the origin is directly below the top of the pile and the z-axis is upward. The cross-section at height z is x^2+y^2=15-z for 0

OpenStudy (math&ing001):

Can you please draw the x,y, and z axis ?

OpenStudy (megannicole51):

i dont have the drawing of it

OpenStudy (agent0smith):

Wouldn't it just be when z=0?

OpenStudy (anonymous):

Yea all points along that circle when z=0

OpenStudy (agent0smith):

x^2+y^2=15-z z=0 gives x^2+y^2=15 So it's a circle, of radius sqrt15

OpenStudy (megannicole51):

oh wow i feel stupid! theres other questions that go with this that i need help with but my book is at home so ill post the other questions later tonight! thank you dear! @agent0smith

OpenStudy (megannicole51):

What is the area of the base of the pile? @agent0smith

OpenStudy (agent0smith):

So it's a circle, of radius sqrt15 Use that to find area Meg :P

OpenStudy (megannicole51):

i have no common sense tonight lol thank youuuu

OpenStudy (megannicole51):

imagine the pile sliced horizontally. What is the area, A(z), of the slice at height z? What are its units?

OpenStudy (agent0smith):

Hmm. Will have to think about that.

OpenStudy (megannicole51):

i cant figure it out...did u?:/

OpenStudy (agent0smith):

Oh it's just horizontal circles!

OpenStudy (agent0smith):

x^2+y^2=15-z the radius is sqrt(15-z) So the area is just using that, and a circle's area!

OpenStudy (megannicole51):

i dont get what you mean:/

OpenStudy (agent0smith):

Radius of x^2+y^2=15 is squrt15 right? So the radius of x^2+y^2=15-z is sqrt(15-z)

OpenStudy (megannicole51):

ooooh i see it now!

OpenStudy (agent0smith):

:)

OpenStudy (megannicole51):

so what does it mean by how many units?

OpenStudy (agent0smith):

I knew you would, you smart girl :D

OpenStudy (agent0smith):

They're all in m, so m^2

OpenStudy (megannicole51):

where did m come from? lol

OpenStudy (agent0smith):

Now you know the radius, use it for... What is the area, A(z), of the slice at height z?

OpenStudy (agent0smith):

Area of a circle

OpenStudy (megannicole51):

pi(15-z)

OpenStudy (agent0smith):

:)

OpenStudy (megannicole51):

okay so now if delta(z) is the thickness of the slice at height z, what is represented by A(z) delta(z)?

OpenStudy (megannicole51):

thanks for helping me with this problem btw

OpenStudy (agent0smith):

Anytime!

OpenStudy (agent0smith):

Area times thickness will give what? Think of a cylinder. What does the area of its base times height give you?

OpenStudy (megannicole51):

how thick the cylinder is?

OpenStudy (agent0smith):

No, for any prism or cylinder, area of base times height gives its...?

OpenStudy (agent0smith):

The "thickness" of the slice is the same thing as its height.

OpenStudy (agent0smith):

How do you find the volume of a cylinder?

OpenStudy (megannicole51):

volume.

OpenStudy (agent0smith):

So the area of a slice times its thickness gives the slice's...?

OpenStudy (megannicole51):

volume

OpenStudy (agent0smith):

:)

OpenStudy (megannicole51):

okay ive got one more question for this question lol evaluate and interpret the integral from 0 to 15 of A(z)dz. what are its units?

OpenStudy (agent0smith):

That'd be the volume of the whole hill thingy. in m^3